miabelle1218:

Mrs. Ramiriz is making pies to sell at the local farmer’s market. It costs her $5 to make each pie, plus a one-time cost of $30 for baking supplies. She plans to sell the pies for $12 each. Which equation can be used to find the number of pies she needs to sell to break even?

YECR:

a.5x – 12x = 30 b.5x + 12x = 30 c.5x + 30 = 12x d.12x + 30 = 5x Choose the correct answer and tell me how u solved it

YECR:

This Question is old though but if you need help my question is up above ^